OpenStudy (anonymous):

Find the derivative of f(x) = 8/x at x = -1. the answer choices are: 4,0,8,-8

OpenStudy (anonymous):

Since I can't just tell you what the answer is, I will tell you what it is not. It is not 4, o, or 8.

OpenStudy (anonymous):

0*

OpenStudy (anonymous):

d/dx (8/x) = 8 x d/dx(1/x) = 8/x^2 substitute the value of x we get 8/-1 = -8

OpenStudy (anonymous):

thanks! @Niall , and ok so to find the derivative, you just plug in x? @paul1231

OpenStudy (anonymous):

Could you help me with one more? @Niall @paul1231 Find the derivative of f(x) = -6/x at x = 12. Answer Choices: 24,1/2,1/24,2

OpenStudy (anonymous):

what is the derivative of 1/x?

OpenStudy (anonymous):

im not sure... how can i find it? @paul1231

OpenStudy (anonymous):

do you know the derivative of x?

OpenStudy (anonymous):

no...i really have no clue how to do this, im not really sure what the 'derivative' is @paul1231

OpenStudy (anonymous):

refer to that and tell me what the derivative of 2x is

OpenStudy (anonymous):

d/dx (2x)= 2x/2x^2? @paul1231

OpenStudy (anonymous):

2f'(2x) = x^2 f'(2x) = x^2 / 2 Let u = 2x. Then x = u/2, so: f'(u) = (u/2)^2 / 2 = u^2 / 8 Thus: f'(x) = x^2 / 8

OpenStudy (anonymous):

d/dx(2x) is 2. DAMNIT